# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@ 5.4 Evidence 2
The main project behind 5.1 is to discover new Hodge cycles for hypersurfaces by deformation. Once such Hodge cycles are discovered, there is an Artinian Gorestein ring attached to such Hodge cycles which contains some partial data of the defining ideal of the underlying algebraic cycle (if the Hodge conjecture is true), see [Voi89, Otw03, MV21]. In the case of lowest codimension for a Hodge locus, this is actually enough to construct the algebraic cycle (in this case a linear cycle) from the topological data of a Hodge cycle, see [Voi89] for $n=2$ and [VL22a] for arbitrary $n$ but near the Fermat variety, and [MS21]. It turns out that in the case of surfaces ($n=2$) the next minimal codimension for Hodge loci (also called Noether-Lefschetz loci) is achieved by surfaces containing a conic, see [Voi89, Voi90] . Therefore, it is expected that components of Hodge loci of low codimension parametrize hypersurfaces with rather simple algebraic cycles. In our case, it turns out that ${\rm dim}(V_{Z})$ grows like the minimal codimension for Hodge loci. This is as follows. A formula for the dimension of $V_{Z}$ for arbitrary $m$ in terms of binomials can be found in [Mov21a, Proposicion 17.9]:
(59) ${\rm codim}(V_{Z})=2{\sf C}_{1^{\frac{n}{2}+1},(d-1)^{\frac{n}{2}+1}}-{\sf C}_{1^{n-m+1},(d-1)^{m+1}}.$
@@
(60) ${\sf C}_{\underline{a}}=\binom{n+1+d}{n+1}-\sum_{k=1}^{2s}(-1)^{k-1}\sum_{a_{i_{1}}+a_{i_{2}}+\cdots+a_{i_{k}}\leq d}\binom{n+1+d-a_{i_{1}}-a_{i_{2}}-\cdots-a_{i_{k}}}{n+1}$
@@ 5.5 Evidence 3
There is a very tiny evidence that the Hodge cycle in 5.1 might be a counterexample to the Hodge conjectures. All the author’s attempts to produce new components of Hodge loci with the same codimension as of $V_{[Z]}$ has failed. This is summarized in [Mov21a, Table 19.5] which we explain it in this section.
@@
For cubic fourfolds, Hodge loci is a union of codimension one irreducible subvarieties ${\cal C}_{D},\ D\equiv_{6}0,2,D\geq 8$ of ${\sf T}$, see [Has00]. Here, $D$ is the discriminant of the saturated lattice generated by $[Z]$ and the polarization $[Z_{\infty}]=[\mathbb{P}^{3}\cap X]$ in $H_{4}(X,\mathbb{Z})$ (in [Has00] notation $[Z_{\infty}]=h^{2}$), where $Z$ is an algebraic cycle $Z\subset X,\ X\in{\cal C}_{D}$ whose homology class together $[Z_{\infty}]$ form a rank two lattice. The loci of cubic fourfolds containing a plane $\mathbb{P}^{2}$ is ${\cal C}_{8}$. It turns out that the generalized $\mathbb{P}^{2}$ is just the linear cycle $\mathbb{P}^{\frac{n}{2}}$ and the space of cubic $n$-folds containing a linear cycle has the smallest possible codimension. These codimesnions are listed under $L$ in Table 1. The loci of cubic fourfolds containing a cubic ruled surface/cubic scroll is ${\cal C}_{12}$. The codimension of the space of cubic $n$-folds containing a generalized cubic scroll is listed in $CS$ in Table 1. Under $M$ we have listed the codimension of our Hodge loci in 5.1. Next comes, ${\cal C}_{14}$ and ${\cal C}_{20}$ for cubic $n$-folds. The loci ${\cal C}_{14}$ parametrizes cubic fourfolds with a quartic scroll. For generalized quartic scroll we get codimensions under $QS$. The loci of cubic fourfolds with a Veronese surface is ${\cal C}_{20}$ and for generalized Veronese we get the codiemsnions under $V$. One gets the impression that as $D$ increases the codimension of any possible generalization of ${\cal C}_{D}$ for cubic hypersurfaces of dimensions $n$ gets near to the maximal codimension, and so, far away from the codimension in 5.1.
@@
Table 1: Codimensions of the components of the Hodge/special loci for cubic hypersurfaces.
@@ 5.6 Artinian Gorenstein ideals attached to Hodge cycles
In order to constuct an algebraic cycle $Z$ from its topological class we must compute its ideal $I_{Z}$ which might be a complicated task. However, we may aim to compute at least one element $g$ of $I_{Z}$ which is not in the ideal $I_{X}$ of the ambient space $X$. In the case of surfaces $X\subset\mathbb{P}^{3}$ this is actually almost the whole task, as we do the intersection $X\cap\mathbb{P}\{g=0\}$, and the only possiblity for $Z$ comes from the irreducible components of this intersection. In general this is as difficult as the original job, and a precise formulation of this has been done in [Tho05]. The linear part of the Artinian-Gorenstein ideal of a Hodge cycle of a hypersurface seems to be part of the defining ideal of the underlying algebraic cycle, and in this section we aim to explain this.
@@
Let ${Z_{\infty}}$ be the intersection of a linear $\mathbb{P}^{\frac{n}{2}+1}$ with $X$ and $[{Z_{\infty}}]\in H_{n}(X,\mathbb{Z})$ be the induced element in homology (the polarization). We have $I([Z])=\mathbb{C}[x]$ and for an arbitrary Hodge cycle $\delta$, $I(\delta)$ depends only on the equivalence class of $\delta\in H_{n}(X,\mathbb{Z})/\mathbb{Z}[{Z_{\infty}}]$. The main purpose of the present section is to investigate the following:
@@
If the Hodge conjecture is true then 5.5 says that the linear polynomial $g$ is in the defining ideal of an algebraic cycle $Z$ such that $\delta=[Z]$. We have the following statement which is stronger than the converse to 5.5. Let $\delta=[Z]\in H_{n}(X,\mathbb{Z})$ be an algebraic cycle. Then the defining ideal of $Z$ is inside $I_{\delta}$. The proof is the same as [MV21, Proposition 11.3].
If we take a basis $g_{1},g_{2},\cdots,g_{k}$ of $(I_{\delta})_{1}$ and apply the above conjecture for $g=\sum_{i=1}^{k}t_{i}g_{i}$ with arbitrary $t_{i}\in\mathbb{C}$ then we may conclude that $\delta$ is supported in $\mathbb{P}\{(I_{\delta})_{1}=0\}\cap X$. A rigorous argument for this is needed, but it does not seem to be difficult. In particular, ${\rm dim}_{\mathbb{C}}(I_{\delta})_{1}\leq\frac{n}{2}+1$. For $X$ the Fermat variety this consequence is easy and it can be reduced to an elementary problem as [Mov21a, Problem 21.3]. 5.5 is mainly inspired by the following conjecture for which we have more evidences.
@@
For $d\not=3,4,6$, $X_{0}$ the Fermat variety and without the smoothness condition this theorem is proved in [VL22a, Theorem 1.2]. For $d=3,4,6$ smoothness is necessary as in [DV21] the authors have described many non-smooth components for which the theorem is not true.
@@
If $Y\subset\mathbb{P}^{n}:=\mathbb{P}\{g=0\}$ is smooth then by Lefschetz’ hyperplane section theorem $H_{n}(Y,\mathbb{Z})\cong H_{n}(\mathbb{P}^{n},\mathbb{Z})$ and the latter is generated by any $\mathbb{P}^{\frac{n}{2}}\subset\mathbb{P}^{n}$. From another side if we take any $\mathbb{P}^{\frac{n}{2}+1}\subset\mathbb{P}^{n}\subset\mathbb{P}^{n+1}$ we have ${Z_{\infty}}\subset Y\subset\mathbb{P}^{n}$, and $[{Z_{\infty}}]=d[\mathbb{P}^{\frac{n}{2}}]$ in $H_{n}(\mathbb{P}^{n},\mathbb{Z})$. This implies that a $d$ multiple of the generator of $H_{n}(Y,\mathbb{Z})$ is $[{Z_{\infty}}]$, and so $\delta$ must be a torsion in $H_{n}(X,\mathbb{Z})/\mathbb{Z}[{Z_{\infty}}]$. ∎
@@ 5.7 Singular cubic hypersurfaces
It follows that the number of $\mathbb{P}^{4}$’s in a generic cubic tenfold is finite. It turns out that such a number is $1812646836$, see [HK22]. For $n=10$ and $k=\frac{n}{2}-2=3$ we have $\delta=8$, that is, the variety of $\mathbb{P}^{3}$’s inside a generic cubic tenfold is of dimension $8$. Next, we focus on singular cubic hypersurfaces.
@@
A singular cubic hypersurface $X\subset\mathbb{P}^{n+1}$ is either a cone over another cubic hypersurface of dimension $n-1$ or it is birational to $\mathbb{P}^{n}$.
@@
If for all point $q\in X$ the line passing through $p$ and $q$ lies in $X$ then the image of $\alpha$ is the point $p$. In this case $X$ is a cone over another cubic hypersurface of dimension $n-1$ and $p$ is the vertex of the cone. Let us assume that this is not the case. Then $\alpha$ is a birational map between $\mathbb{P}^{n}_{p}$ and $X$. ∎
It is useful to rewrite the above proof in a coordinate system $[x_{0}:x_{1}:\cdots:x_{n+1}]$. We take the affine chart $x=(x_{1},x_{2},\ldots,x_{n+1})\in\mathbb{C}^{n}$ given by $x_{0}=1$ and assume that the singularity $p$ is at the origin $0\in\mathbb{C}^{n+1}$. The hypersurface $X$ is given by $f=x_{0}f_{2}-f_{3}$, where $f_{i}$’ are homogenuous polynomials of degree $i$ in $x$. If $f_{2}=0$ then $X$ is a cone over the cubic hypersurface $\mathbb{P}\{f_{3}=0\}\subset\mathbb{P}^{n}$. Otherwise, we have the birational map
@@
(61) ${\rm Sing}\{f_{2}=0\}\cap{\rm Sing}\{f_{3}=0\}\subset{\rm Sing}(X)\cap\mathbb{C}^{n+1}\subset\{f_{2}=0\}\cap\{f_{3}=0\}$
(62) ${\rm Sing}(X)\cap\mathbb{P}^{n}_{\infty}={\rm Sing}\mathbb{P}\{f_{3}=0\}\cap\mathbb{P}\{f_{2}=0\}.$
Moreover, any line between $0\in\mathbb{C}^{n+1}$ and $p\in{\rm Sing}(X)\cap\mathbb{C}^{n+1}$ either lies in ${\rm Sing}(X)$ for which $p\in{\rm Sing}(f_{2}=0)\cap{\rm Sing}(f_{3}=0)$ or it intersects ${\rm Sing}(X)$ only at 00 and $p$.
@@
The variety $X$ is given by $x_{0}f_{2}(x)-f_{3}(x)=0$ and hence ${\rm Sing}(X)$ is given by $x_{0}f_{2}(x)-f_{3}(x)=f_{2}=x_{0}\frac{\partial f_{2}}{\partial x_{i}}-\frac{\partial f_{3}}{\partial x_{i}}=0,\ i=1,2,\ldots,n+1$. The inclusions (61) and (62) are immediate. ∎
@@ 5.8 Computing Artinian Gorenstein ring over formal power series
The hypersurface $X_{t},\ \ t\in V_{[Z]}\backslash V_{Z}$ is not given explicitly, as its existence is conjectural. Therefore, it might be difficult to study its Artinian Gorenstein ring. However, as we can write the Taylor series of the periods of $X_{t},t\in({\sf T},0)$ explicitely, see [Mov21a, Sections 13.9, 13.10, 18.5] we might try to study such rings over, not only over $\mathbb{C}$, but also over formal power series. In this section we explain this idea.
In [Mov21a, Section 19.3], we have taken a parameter space which is transversal to $V_{Z}$ at 00 and it has the complimentary dimension. Therefore, it intersects $V_{Z}$ only at 00. From now on we use $V_{Z}$ and $V_{[Z]}$ for this new parameter space, and hence by our construction $V_{Z}=\{0\}$. 5.1 is equivalent to the follwing: The Hodge locus $V_{[Z]}$ is a smooth curve (${\rm dim}(V_{[Z]})=1$). We note that 5.3 is proved first for this new parameter space. In particular, this implies that the new parameter space is also transversal to $T_{0}V_{[Z]}$.
For a smooth hypersurface defined over the ring ${\cal O}_{{\sf T},0}$ of holomorphic functions in a neighborhood of 00, and a continuous family of cycles $\delta=\delta_{t}\in H_{n}(X_{t},\mathbb{Z})/\mathbb{Z}[{Z_{\infty}}],\ t\in({\sf T},0)$, the Hodge locus $V_{\delta}$ is given by the zero locus of an ideal ${\cal I}(\delta)\subset{\cal O}_{{\sf T},0}$.
@@
(63) $I(\delta)_{a}:=\left\{Q\in{\cal O}_{{\sf T},0}[x]_{a}\Bigg{|}{\int}_{\delta_{t}}{\rm res}\left(\frac{QP\Omega}{F_{t}^{\frac{n}{2}+1}}\right)\in{\cal I}(\delta),\ \ \forall P\in\mathbb{C}[x]_{\sigma-a}\right\}.$
We define the Artinian Gorenstein algebra of the Hodge locus as $R(\delta):={\cal O}_{{\sf T},0}[x]/I(\delta)$. By definition $I(\delta)_{m}={\cal O}_{{\sf T},0}[x]_{m}$ for all $m\geq\sigma+1$ and so $R(\delta)_{a}=0$.
Note that we actually need that the integral in (63) vanishes identically over ${\rm Zero}({\cal I}(\delta))$. Since $\cal I(\delta)$ might not be reduced, these two definitions might not be equivalent. Since in 5.1 we expect that $V_{[Z]}$ is smooth, these two definitions are the same. In a similar way we can replace ${\cal O}_{{\sf T},0}$ with the ring $\check{\cal O}_{{\sf T},0}$ of formal power series, and in particular, with the truncated rings ${\cal O}_{{\sf T},0}^{N}:={\cal O}_{{\sf T},0}/m_{{\sf T},0}^{N+1}\cong\check{\cal O}_{{\sf T},0}/\check{m}_{{\sf T},0}^{N+1}$.
@@
It seems quite possible to prove this conjecture using [Voi88, Section 3] and [Otw02, Theorem 3, Proposition 6]. In these reference the authors prove that if a Hodge locus $V_{\delta}$ has minimal codimension then ${\rm dim}I(\delta)=\frac{n}{2}+1$. Note that the codimension of our Hodge locus as a function in $n$ grows as the minimal codimension for a Hodge loci, see Section 5.4. Despite this, we want to get some evidence for 5.7. The main goal of this section is to explain the computer code which verifies the following statement: The linear part $I^{N}(\delta)_{1}$ of $I^{N}(\delta)$ is not zero for $(n,N)=(6,7),(8,3)$. For the computer code for experimental verification of 5.7 see the tex file of the present text in arxiv or the author’s webpage.
@@
For the Hodge cycle in 5.1 we want to compute $I(\delta)_{1}$ which is equivalent to compute the kernel of $A_{t}$ modulo ${\cal I}(\delta)$ from the left, that is $1\times a_{1}$ vectors $v$ with $vA_{t}=0$ modulo ${\cal I}(\delta)$. At first step we aim to compute the rank of $A_{t}$. Let $\mu$ be the rank of $A_{t}$ over ${\cal O}_{{\sf T},0}/{\cal I}(\delta)$. This means that the determinant of all $(\mu+1)\times(\mu+1)$ minors of $A_{t}$ are in the ideal modulo ${\cal I}(\delta)$, but there is a $\mu\times\mu$ minor whose determinant is not in ${\cal I}(\delta)$. Recall that ${\cal I}(\delta)$ is conjecturally reduced! These statements can be experimented by computer after truncating the entries of $A_{t}$.

### v2

@@ 5.4 Evidence 2
The main project behind 5.1 is to discover new Hodge cycles for hypersurfaces by deformation. Once such Hodge cycles are discovered, there is an Artinian Gorestein ring attached to such Hodge cycles which contains some partial data of the defining ideal of the underlying algebraic cycle (if the Hodge conjecture is true), see [voisin89, Otwinowska2003, ho13-Roberto]. In the case of lowest codimension for a Hodge locus, this is actually enough to construct the algebraic cycle (in this case a linear cycle) from the topological data of a Hodge cycle, see [voisin89] for $n=2$ and [Villaflor2022] for arbitrary $n$ but near the Fermat variety, and [EmreHossein2018]. It turns out that in the case of surfaces ($n=2$) the next minimal codimension for Hodge loci (also called Noether-Lefschetz loci) is achieved by surfaces containing a conic, see [voisin89, voisin90] . Therefore, it is expected that components of Hodge loci of low codimension parametrize hypersurfaces with rather simple algebraic cycles. In our case, it turns out that ${\rm dim}(V_{Z})$ grows like the minimal codimension for Hodge loci. This is as follows. A formula for the dimension of $V_{Z}$ for arbitrary $m$ in terms of binomials can be found in [ho13, Proposicion 17.9]:
(5.6) ${\rm codim}(V_{Z})=2{\sf C}_{1^{\frac{n}{2}+1},(d-1)^{\frac{n}{2}+1}}-{\sf C}_{1^{n-m+1},(d-1)^{m+1}}.$
@@
(5.7) ${\sf C}_{\underline{a}}=\binom{n+1+d}{n+1}-\sum_{k=1}^{2s}(-1)^{k-1}\sum_{a_{i_{1}}+a_{i_{2}}+\cdots+a_{i_{k}}\leq d}\binom{n+1+d-a_{i_{1}}-a_{i_{2}}-\cdots-a_{i_{k}}}{n+1}$
@@ 5.5 Evidence 3
There is a very tiny evidence that the Hodge cycle in 5.1 might be a counterexample to the Hodge conjectures. All the author’s attempts to produce new components of Hodge loci with the same codimension as of $V_{[Z]}$ has failed. This is summarized in [ho13, Table 19.5] which we explain it in this section.
@@
For cubic fourfolds, Hodge loci is a union of codimension one irreducible subvarieties ${\mathcal{C}}_{D},\ D\equiv_{6}0,2,D\geq 8$ of $\sf T$}, see [Hassett2000]. Here, $D$ is the discriminant of the saturated lattice generated by $[Z]$ and the polarization $Z_{\infty}]=[\mathbb{P}^{3}\cap X]$ in $H_{4}(X,\mathbb{Z})$ (in [Hassett2000] notation $Z_{\infty}]=h^{2}$), where $Z$ is an algebraic cycle $Z\subset X,\ X\in{\mathcal{C}}_{D}$ whose homology class together $Z_{\infty}]$ form a rank two lattice. The loci of cubic fourfolds containing a plane $\mathbb{P}^{2}$ is ${\mathcal{C}}_{8}$. It turns out that the generalized $\mathbb{P}^{2}$ is just the linear cycle $\mathbb{P}^{\frac{n}{2}}$ and the space of cubic $n$-folds containing a linear cycle has the smallest possible codimension. These codimesnions are listed under $L$ in Table 5.1. The loci of cubic fourfolds containing a cubic ruled surface/cubic scroll is ${\mathcal{C}}_{12}$. The codimension of the space of cubic $n$-folds containing a generalized cubic scroll is listed in $CS$ in Table 5.1. Under $M$ we have listed the codimension of our Hodge loci in 5.1. Next comes, ${\mathcal{C}}_{14}$ and ${\mathcal{C}}_{20}$ for cubic $n$-folds. The loci ${\mathcal{C}}_{14}$ parametrizes cubic fourfolds with a quartic scroll. For generalized quartic scroll we get codimensions under $QS$. The loci of cubic fourfolds with a Veronese surface is ${\mathcal{C}}_{20}$ and for generalized Veronese we get the codiemsnions under $V$. One gets the impression that as $D$ increases the codimension of any possible generalization of ${\mathcal{C}}_{D}$ for cubic hypersurfaces of dimensions $n$ gets near to the maximal codimension, and so, far away from the codimension in 5.1.
@@ 88 $120$ $10,45$ 10 16 19 23 25 $0,0,0,45,253,45,0,0,0$
Table 5.1: Codimensions of the components of the Hodge/special loci for cubic hypersurfaces.
@@ 5.6 Artinian Gorenstein ideals attached to Hodge cycles
In order to constuct an algebraic cycle $Z$ from its topological class we must compute its ideal $I_{Z}$ which might be a complicated task. However, we may aim to compute at least one element $g$ of $I_{Z}$ which is not in the ideal $I_{X}$ of the ambient space $X$. In the case of surfaces $X\subset\mathbb{P}^{3}$ this is actually almost the whole task, as we do the intersection $X\cap\mathbb{P}\{g=0\}$, and the only possiblity for $Z$ comes from the irreducible components of this intersection. In general this is as difficult as the original job, and a precise formulation of this has been done in [Thomas2005]. The linear part of the Artinian-Gorenstein ideal of a Hodge cycle of a hypersurface seems to be part of the defining ideal of the underlying algebraic cycle, and in this section we aim to explain this.
@@
Let $Z_{\infty}$} be the intersection of a linear $\mathbb{P}^{\frac{n}{2}+1}$ with $X$ and ${Z_{\infty}}]\in H_{n}(X,\mathbb{Z})$ be the induced element in homology (the polarization). We have $I([Z])=\mathbb{C}[x]$ and for an arbitrary Hodge cycle $\delta$, $I(\delta)$ depends only on the equivalence class of $\delta\in H_{n}(X,\mathbb{Z})/\mathbb{Z}[{Z_{\infty}}]$. The main purpose of the present section is to investigate the following:
@@
If the Hodge conjecture is true then 5.5 says that the linear polynomial $g$ is in the defining ideal of an algebraic cycle $Z$ such that $\delta=[Z]$. We have the following statement which is stronger than the converse to 5.5. Let $\delta=[Z]\in H_{n}(X,\mathbb{Z})$ be an algebraic cycle. Then the defining ideal of $Z$ is inside $I_{\delta}$. The proof is the same as [ho13-Roberto, Proposition 11.3].
If we take a basis $g_{1},g_{2},\cdots,g_{k}$ of $(I_{\delta})_{1}$ and apply the above conjecture for $g=\sum_{i=1}^{k}t_{i}g_{i}$ with arbitrary $t_{i}\in\mathbb{C}$ then we may conclude that $\delta$ is supported in $\mathbb{P}\{(I_{\delta})_{1}=0\}\cap X$. A rigorous argument for this is needed, but it does not seem to be difficult. In particular, ${\rm dim}_{\mathbb{C}}(I_{\delta})_{1}\leq\frac{n}{2}+1$. For $X$ the Fermat variety this consequence is easy and it can be reduced to an elementary problem as [ho13, Problem 21.3]. 5.5 is mainly inspired by the following conjecture for which we have more evidences.
@@
For $d\not=3,4,6$, $X_{0}$ the Fermat variety and without the smoothness condition this theorem is proved in [Villaflor2022, Theorem 1.2]. For $d=3,4,6$ smoothness is necessary as in [DuqueVillaflor] the authors have described many non-smooth components for which the theorem is not true.
@@
If $Y\subset\mathbb{P}^{n}:=\mathbb{P}\{g=0\}$ is smooth then by Lefschetz’ hyperplane section theorem $H_{n}(Y,\mathbb{Z})\cong H_{n}(\mathbb{P}^{n},\mathbb{Z})$ and the latter is generated by any $\mathbb{P}^{\frac{n}{2}}\subset\mathbb{P}^{n}$. From another side if we take any $\mathbb{P}^{\frac{n}{2}+1}\subset\mathbb{P}^{n}\subset\mathbb{P}^{n+1}$ we have ${Z_{\infty}}\subset Y\subset\mathbb{P}^{n}$, and ${Z_{\infty}}]=d[\mathbb{P}^{\frac{n}{2}}]$ in $H_{n}(\mathbb{P}^{n},\mathbb{Z})$. This implies that a $d$ multiple of the generator of $H_{n}(Y,\mathbb{Z})$ is $Z_{\infty}$}], and so $\delta$ must be a torsion in $H_{n}(X,\mathbb{Z})/\mathbb{Z}[{Z_{\infty}}]$. ∎
@@ 5.7 Singular cubic hypersurfaces
It follows that the number of $\mathbb{P}^{4}$’s in a generic cubic tenfold is finite. It turns out that such a number is $1812646836$, see [HashimotoKadets]. For $n=10$ and $k=\frac{n}{2}-2=3$ we have $\delta=8$, that is, the variety of $\mathbb{P}^{3}$’s inside a generic cubic tenfold is of dimension $8$. Next, we focus on singular cubic hypersurfaces.
@@
A singular cubic hypersurface $X\subset\mathbb{P}^{n+1}$ is either a cone over another cubic hypersurface of dimension n1n-1 or it is birational to $\mathbb{P}^{n}$.
@@
If for all point $q\in X$ the line passing through $p$ and $q$ lies in $X$ then the image of $\alpha$ is the point $p$. In this case $X$ is a cone over another cubic hypersurface of dimension n1n-1 and $p$ is the vertex of the cone. Let us assume that this is not the case. Then $\alpha$ is a birational map between $\mathbb{P}^{n}_{p}$ and $X$. ∎
It is useful to rewrite the above proof in a coordinate system $x_{0}:x_{1}:\cdots:x_{n+1}]$. We take the affine chart $x=(x_{1},x_{2},\ldots,x_{n+1})\in\mathbb{C}^{n}$ given by $x_{0}=1$ and assume that the singularity $p$ is at the origin $0\in\mathbb{C}^{n+1}$. The hypersurface $X$ is given by $f=x_{0}f_{2}-f_{3}$, where $f_{i}$’ are homogenuous polynomials of degree $i$ in $x$. If $f_{2}=0$ then $X$ is a cone over the cubic hypersurface $\mathbb{P}\{f_{3}=0\}\subset\mathbb{P}^{n}$. Otherwise, we have the birational map
@@
(5.8) ${\rm Sing}\{f_{2}=0\}\cap{\rm Sing}\{f_{3}=0\}\subset{\rm Sing}(X)\cap\mathbb{C}^{n+1}\subset\{f_{2}=0\}\cap\{f_{3}=0\}$
(5.9) ${\rm Sing}(X)\cap\mathbb{P}^{n}_{\infty}={\rm Sing}\mathbb{P}\{f_{3}=0\}\cap\mathbb{P}\{f_{2}=0\}.$
Moreover, any line between $0\in\mathbb{C}^{n+1}$ and $p\in{\rm Sing}(X)\cap\mathbb{C}^{n+1}$ either lies in ${\rm Sing}(X)$ for which $p\in{\rm Sing}(f_{2}=0)\cap{\rm Sing}(f_{3}=0)$ or it intersects ${\rm Sing}(X)$ only at 0 and $p$.
@@
The variety $X$ is given by $x_{0}f_{2}(x)-f_{3}(x)=0$ and hence ${\rm Sing}(X)$ is given by $x_{0}f_{2}(x)-f_{3}(x)=f_{2}=x_{0}\frac{\partial f_{2}}{\partial x_{i}}-\frac{\partial f_{3}}{\partial x_{i}}=0,\ i=1,2,\ldots,n+1$. The inclusions (5.8) and (5.9) are immediate. ∎
@@ 5.8 Computing Artinian Gorenstein ring over formal power series
The hypersurface $X_{t},\ \ t\in V_{[Z]}\backslash V_{Z}$ is not given explicitly, as its existence is conjectural. Therefore, it might be difficult to study its Artinian Gorenstein ring. However, as we can write the Taylor series of the periods of $X_{t},t\in({\sf T},0)$ explicitely, see [ho13, Sections 13.9, 13.10, 18.5] we might try to study such rings over, not only over $\mathbb{C}$, but also over formal power series. In this section we explain this idea.
In [ho13, Section 19.3], we have taken a parameter space which is transversal to $V_{Z}$ at 0 and it has the complimentary dimension. Therefore, it intersects $V_{Z}$ only at 0. From now on we use $V_{Z}$ and $V_{[Z]}$ for this new parameter space, and hence by our construction $V_{Z}=\{0\}$. 5.1 is equivalent to the follwing: The Hodge locus $V_{[Z]}$ is a smooth curve (${\rm dim}(V_{[Z]})=1$). We note that 5.3 is proved first for this new parameter space. In particular, this implies that the new parameter space is also transversal to $T_{0}V_{[Z]}$.
For a smooth hypersurface defined over the ring ${\mathcal{O}}_{{\sf T},0}$ of holomorphic functions in a neighborhood of 0, and a continuous family of cycles $\delta=\delta_{t}\in H_{n}(X_{t},\mathbb{Z})/\mathbb{Z}[{Z_{\infty}}],\ t\in({\sf T},0)$, the Hodge locus $V_{\delta}$ is given by the zero locus of an ideal ${\mathcal{I}}(\delta)\subset{\mathcal{O}}_{{\sf T},0}$.
@@
(5.10) $I(\delta)_{a}:=\left\{Q\in{\mathcal{O}}_{{\sf T},0}[x]_{a}\Bigg|{\int}_{\delta_{t}}{\rm res}\left(\frac{QP\Omega}{F_{t}^{\frac{n}{2}+1}}\right)\in{\mathcal{I}}(\delta),\ \ \forall P\in\mathbb{C}[x]_{\sigma-a}\right\}.$
We define the Artinian Gorenstein algebra of the Hodge locus as $R(\delta):={\mathcal{O}}_{{\sf T},0}[x]/I(\delta)$. By definition $I(\delta)_{m}={\mathcal{O}}_{{\sf T},0}[x]_{m}$ for all $m\geq\sigma+1$ and so $R(\delta)_{a}=0$.
Note that we actually need that the integral in (5.10) vanishes identically over ${\rm Zero}({\mathcal{I}}(\delta))$. Since $\mathcal{I}(\delta)$ might not be reduced, these two definitions might not be equivalent. Since in 5.1 we expect that $V_{[Z]}$ is smooth, these two definitions are the same. In a similar way we can replace ${\mathcal{O}}_{{\sf T},0}$ with the ring $\check{\mathcal{O}}_{{\sf T},0}$ of formal power series, and in particular, with the truncated rings ${\mathcal{O}}_{{\sf T},0}^{N}:={\mathcal{O}}_{{\sf T},0}/m_{{\sf T},0}^{N+1}\cong\check{\mathcal{O}}_{{\sf T},0}/\check{m}_{{\sf T},0}^{N+1}$.
@@
It seems quite possible to prove this conjecture using [voisin1988, Section 3] and [Otwinowska2002, Theorem 3, Proposition 6]. In these reference the authors prove that if a Hodge locus $V_{\delta}$ has minimal codimension then ${\rm dim}I(\delta)=\frac{n}{2}+1$. Note that the codimension of our Hodge locus as a function in $n$ grows as the minimal codimension for a Hodge loci, see Section 5.4. Despite this, we want to get some evidence for 5.7. The main goal of this section is to explain the computer code which verifies the following statement: The linear part $I^{N}(\delta)_{1}$ of $I^{N}(\delta)$ is not zero for $(n,N)=(6,7),(8,3)$. For the computer code for experimental verification of 5.7 see the tex file of the present text in arxiv or the author’s webpage.
@@
For the Hodge cycle in 5.1 we want to compute $I(\delta)_{1}$ which is equivalent to compute the kernel of $A_{t}$ modulo ${\mathcal{I}}(\delta)$ from the left, that is $1\times a_{1}$ vectors $v$ with $vA_{t}=0$ modulo ${\mathcal{I}}(\delta)$. At first step we aim to compute the rank of $A_{t}$. Let $\mu$ be the rank of $A_{t}$ over ${\mathcal{O}}_{{\sf T},0}/{\mathcal{I}}(\delta)$. This means that the determinant of all $(\mu+1)\times(\mu+1)$ minors of $A_{t}$ are in the ideal modulo ${\mathcal{I}}(\delta)$, but there is a $\mu\times\mu$ minor whose determinant is not in ${\mathcal{I}}(\delta)$. Recall that ${\mathcal{I}}(\delta)$ is conjecturally reduced! These statements can be experimented by computer after truncating the entries of $A_{t}$.